The metal bar is AB =10mm, DA=18mm, BC =220cm and the resistivity of the metal is 0,012 micro ohms. what is its tip to tip resistance


1
Expert's answer
2021-06-08T06:53:02-0400

Resistance of a conductor of uniform section "S" is "R =\frac{\rho l}{S}", where "\rho" is electrical resistivity, "l" is length.

The area of the cross section of the bar is "S = 10 \cdot 10^{-3} m \cdot 18 \cdot 10^{-3} m = 1.8 \cdot 10^{-4} m^2", and the length is "l = 2.2 m". Hence, the resistance is:

"R = \frac{0.012 \cdot 10^{-6} \Omega \cdot m \ 2.2 m\cdot }{1.8 \cdot 10^{-4} m^2} \approx 1.5 \cdot 10^{-4} \Omega".